# Comparative study of magnetic exchange parameters and magnon dispersions in NiO and MnO from first principles

### v1

@@ RESULTS
NiO and MnO are prototypical insulating transition-metal monoxides, widely used to benchmark and compare computational methods for describing structural, electronic, magnetic, and spectral properties. At high temperatures, both materials crystallize in the rocksalt-type structure and adopt a paramagnetic phase. Upon cooling below their respective Néel temperatures, 524 K for NiO 110 and 120 K for MnO 111, they undergo a transition to a type II antiferromagnetic (AFM) phase. This magnetic ordering is accompanied by a small rhombohedral distortion along the [111] direction of the face-centered cubic (fcc) lattice. In the AFM phase, ferromagnetic (111) planes are stacked antiferromagnetically along the [111] direction, as illustrated in Fig. 1.
@@
Due to the small rhombohedral distortion of the lattice at low temperatures, certain exchange interactions become inequivalent for parallel and antiparallel spin orientations. In particular, the nearest-neighbor exchange interaction $J_{1}$ splits into two distinct parameters, $J_{1}^{-}$ and $J_{1}^{+}$ – a notation introduced by Lines and Jones 112 – to distinguish between interactions involving parallel and antiparallel spin pairs, respectively. This splitting originates from magnetoelastic coupling induced by the distortion. In the absence of such structural distortions, these parameters are equal, i.e., $J_{1}^{-}=J_{1}^{+}$. In contrast, the next-nearest-neighbor exchange interaction $J_{2}$ occurs only between antiparallel spin pairs, and thus a single parameter is sufficient to describe it. These exchange interactions are illustrated in Fig. 1; interactions beyond the second-neighbor shell are not shown. A detailed analysis of the physical origin and behavior of exchange interactions in these systems can be found, for example, in Refs. 24, 25.
@@
First-principles calculations of exchange parameters using the direct MFT and $\Delta E$ methods
@@
As discussed, we first compute magnetic exchange parameters using two direct first-principles approaches: MFT and $\Delta E$. Both methods are applied consistently on top of the same DFT+$U$ ground state, using identical Hubbard $U$ values obtained from DFPT. The inclusion of $U$ is essential for NiO and MnO, as it corrects self-interaction errors of (semi-)local functionals and thus properly localizes the $d$ electrons. This consistent setup contrasts with previous studies 25, 38, 35, 106, where a single method was used to compute exchange parameters but with varying ground-state descriptions. By fixing the ground state using DFT+$U$, we investigate the accuracy of each method (MFT and $\Delta E$) in predicting the MIPs, similar to the strategy in Ref. 24 (which employed the SIC method for the ground-state description). Moreover, using first-principles $U$ values removes ambiguity in the choice of $U$, avoiding additional spread in the computed MIPs.
@@
Each of the methods considered yields individual exchange parameters $J$, which must be computed for as many neighbor shells as needed to ensure convergence of the magnon dispersions. For both NiO and MnO, we find that including exchange interactions up to the fifth nearest neighbor shell is necessary to achieve stable (i.e., positive-definite) magnon dispersions with convergence within a few meV. The corresponding values are listed in Table 1. While most previous studies 105, 24, 25, 38, 35, 106, 107, 108 limit the analysis to only $J_{1}$ and $J_{2}$ (often neglecting the splitting of $J_{1}$), our results highlight the importance of computing further-neighbor couplings to obtain stable and quantitatively accurate magnon dispersions. We also find that the splitting between parallel and antiparallel spin interactions is relevant not only for the first nearest-neighbor shell ($J_{1}$) but also for the third nearest-neighbor shell ($J_{3}$). All reported $J$ values are expressed using spin quantum numbers $S=1$ for Ni2+ and $S=5/2$ for Mn2+ (see Eq. (1)). Alternative definitions, such as setting $S$ equal to half the calculated magnetic moment, exist in the literature, but suffer from ambiguity in the definition of the magnetic moment. Therefore, the choice of $S$ must be taken into account when comparing the exchange parameters across different computational studies or with experimentally fitted values, as will be discussed below.
@@
Figure 2: Magnon dispersions for (a) NiO and (b) MnO computed using LSWT with Heisenberg exchange parameters from MFT and $\Delta E$ are shown as blue and red solid lines, respectively (see Table 1 and Eq. (1)). The TDDFPT+$U$ magnons dispersions are computed directly from the dynamical spin susceptibility and are shown as green dots. Experimental magnon dispersions, shown as black dots, are taken from inelastic neutron scattering measurements reported in Refs. 113, 114. The green and black dashed lines represent LSWT fits to the TDDFPT+$U$ and experimental data, respectively, based on the effective Heisenberg Hamiltonian of Eq. (2). All theoretical results are based on the same DFT+$U$ ground state.
@@
For NiO, both MFT and $\Delta E$ yield similar values for the dominant antiferromagnetic exchange interaction $J_{2}$, with MFT predicting a value approximately 12% higher than that from $\Delta E$. Consistently with previous computational studies, $J_{2}$ is the largest exchange parameter. In contrast, $J_{1}^{+}$ and $J_{1}^{-}$ are 1–2 orders of magnitude smaller and have opposite sign to $J_{2}$. However, their values differ significantly between methods: MFT predicts $|J_{1}^{+}|<|J_{1}^{-}|$, which favors the stabilization of the observed antiferromagnetic spin structure. In contrast, $\Delta E$ gives $|J_{1}^{+}|>|J_{1}^{-}|$, which destabilizes the magnon spectrum unless further-neighbor interactions are included. For the third to fifth neighbor interactions, both methods predict very small values. MFT yields negligible values, with only $J_{4}$ being about 1.5–2 times smaller than $J_{1}^{+}$ and $J_{1}^{-}$. A similar trend is observed in $\Delta E$, where only $J_{4}$ is sizable. Although these couplings are minor compared to $J_{2}$, they are essential in the $\Delta E$ approach to stabilize magnons, while their impact in MFT is minimal [see Fig. S1(a) in the Supporting Information (SI)].
@@
For MnO, the dominant exchange interactions are $J_{1}^{+}$, $J_{1}^{-}$, and $J_{2}$, all having similar magnitudes and the same (positive) sign within each method, unlike in NiO. However, MFT and $\Delta E$ predict different trends: MFT gives $|J_{1}^{+}|>|J_{2}|>|J_{1}^{-}|$, while $\Delta E$ yields $|J_{2}|>|J_{1}^{+}|>|J_{1}^{-}|$. The inclusion of these three parameters alone is sufficient to stabilize magnons in both approaches. However, their absolute values are consistently larger in MFT than in $\Delta E$ by roughly a factor of two, leading to similarly scaled differences in magnon energies [see Fig. S1(b) in the SI]. Third to fifth nearest-neighbor interactions are 1–2 orders of magnitude smaller than $J_{1}^{+}$, $J_{1}^{-}$, and $J_{2}$. Despite their small magnitude, including these further-neighbor interactions is necessary to achieve convergence of the magnon dispersion. Omitting them leads to energy deviations of up to 4 meV for certain magnon momenta, corresponding to errors of about 15%. Thus, while these interactions are not essential for magnon stability (as they are in NiO) they are important for the quantitative comparison with the experimental data.
Magnon dispersions
@@
We now compare the three computational approaches considered, to predict magnon dispersions in NiO and MnO. The first two are based on LSWT using Heisenberg exchange parameters $J$ obtained from MFT and $\Delta E$, as discussed in the previous section. The third approach computes magnon dispersions directly via TDDFPT+$U$ by evaluating the dynamical spin susceptibility 39. A brief overview of each method is provided in Sec. S3 of the SI. All methods are applied consistently on top of the same DFT+$U$ ground state, enabling a direct comparison focused on the physical approximations inherent to each method in describing magnetic excitations. The computed magnon spectra are benchmarked against inelastic neutron scattering data from Refs. 113, 114.
@@
The primitive magnetic unit cell of NiO and MnO contains two antiferromagnetically aligned magnetic ions due to the AFII spin configuration. In the absence of magnetic anisotropy, this leads to two degenerate magnon modes. However, weak magnetocrystalline anisotropies (arising from spin-orbit coupling) lift this degeneracy near the $\Gamma$ point, resulting in a small magnon gap at $\Gamma$ of less than 5 meV in both materials 112, 113, 114. Although this gap is small in NiO and more pronounced in MnO, relative to the maximum magnon energies in each material, we follow common practice in the literature and neglect magnetocrystalline anisotropy, treating the magnon gap at $\Gamma$ as zero. This simplification is justified because our primary aim is to compare the accuracy of different computational methods in predicting magnon dispersions across the full BZ, rather than focusing on fine features near the $\Gamma$ point.
@@
Figure 2 shows the magnon dispersions computed using the three methods along the $\Gamma$–M and $\Gamma$–X directions in the BZ for rhombohedrally distorted NiO and MnO. TDDFPT+$U$ and LSWT based on $\Delta E$-derived exchange parameters both reproduce the experimental data well, capturing the overall trends and curvatures of the magnon bands. The maximum magnon energies are overestimated on average by about 15 meV (14%) for NiO and 3 meV (17%) for MnO. Between the two methods, LSWT with $\Delta E$-derived parameters shows slightly better agreement with experiment than TDDFPT+$U$. This good agreement of both methods with experiments is remarkable given that both methods are fully first-principles and rely exclusively on ab initio Hubbard $U$ values, without any empirical fitting. In contrast, LSWT using MFT-derived $J$ values, while qualitatively reproducing the general features of the dispersion, significantly overestimates the magnon energies. The deviations reach 46 meV (46%) for NiO and 25 meV (139%) for MnO. Such a large deviation of magnon energies from MFT, especially for MnO, is quite striking when compared to the other two methods.
@@
A key feature of the magnon dispersion in NiO and MnO is the finite magnon energy at the M point in the BZ. This feature arises only when the rhombohedral lattice distortion is taken into account 35, 39, which splits the nearest-neighbor (and some further-neighbor) exchange interaction into two distinct parameters, $J_{1}^{+}$ and $J_{1}^{-}$. If the distortion is neglected, then $J_{1}^{+}=J_{1}^{-}$, and the magnon energy at M becomes zero. In our calculations, which include the rhombohedral distortion, all three methods correctly predict a finite magnon energy at M. However, the values obtained from TDDFPT+$U$ and LSWT based on $\Delta E$-derived exchange parameters agree much better with experiment than those from LSWT based on MFT-derived parameters. This magnon energy at M is relatively small in NiO compared to its maximum magnon energy, but significantly larger in MnO. This difference reflects the relative magnitudes of the exchange interactions: in NiO, $J_{2}$ dominates and exceeds $J_{1}^{+}$ and $J_{1}^{-}$ by 1–2 orders of magnitude, whereas in MnO, all three parameters are of similar magnitude (see Table 1).
@@
Figure 3: Effective exchange parameters for NiO within the minimal Heisenberg model [Eq. (2)]: (a) $\mathcal{J}_{1}^{+}$, (b) $\mathcal{J}_{1}^{-}$, and (c) $\mathcal{J}_{2}$, all in meV, extracted by fitting magnon dispersions using the analytical LSWT expression (see Sec. S3.3 in the SI) for four data sets: $(i)$ LSWT dispersions with MFT-derived $J_{1}$$J_{5}$ values, $(ii)$ LSWT dispersions with $\Delta E$-derived $J_{1}$$J_{5}$ values, $(iii)$ TDDFPT+$U$ magnon dispersions from the dynamical spin susceptibility, $(iv)$ Experimental data from Ref. 113. Theoretical results are shown as histograms, while experimental values are indicated by horizontal dashed lines. All theoretical results are based on the same DFT+$U$ ground state.
@@
Figure 4: Effective exchange parameters for MnO within the minimal Heisenberg model [Eq. (2)]: (a) $\mathcal{J}_{1}^{+}$, (b) $\mathcal{J}_{1}^{-}$, and (c) $\mathcal{J}_{2}$, all in meV, extracted by fitting magnon dispersions using the analytical LSWT expression (see Sec. S3.3 in the SI) for four data sets: $(i)$ LSWT dispersions with MFT-derived $J_{1}$$J_{5}$ values, $(ii)$ LSWT dispersions with $\Delta E$-derived $J_{1}$$J_{5}$ values, $(iii)$ TDDFPT+$U$ magnon dispersions from the dynamical spin susceptibility, $(iv)$ Experimental data from Ref. 114. Theoretical results are shown as histograms, while experimental values are indicated by horizontal dashed lines. All theoretical results are based on the same DFT+$U$ ground state.
Fitting effective exchange parameters of the minimal Heisenberg model

### v2

@@ III.1 Magnetic order and Heisenberg Hamiltonians for NiO and MnO
NiO and MnO are prototypical insulating transition-metal monoxides, widely used to benchmark and compare computational methods for describing structural, electronic, magnetic, and spectral properties. At high temperatures, both materials crystallize in the rocksalt-type structure and adopt a paramagnetic phase. Upon cooling below their respective Néel temperatures, 524 K for NiO [126] and 120 K for MnO [127], they undergo a transition to a type II antiferromagnetic (AFM) phase. This magnetic ordering is accompanied by a small rhombohedral distortion along the [111] direction of the face-centered cubic (fcc) lattice. In the AFM phase, ferromagnetic (111) planes are stacked antiferromagnetically along the [111] direction, as illustrated in Fig. 1.
Figure 1: Face-centered cubic conventional unit cell of NiO and MnO with the AFII magnetic order. Transition-metal elements (Ni in NiO, or Mn in MnO) and O atoms are shown in purple and grey, respectively. Yellow and blue vertical thick arrows centered on atoms indicate the direction of the spin. The ferromagnetic (111) planes are depicted in transparent blue and yellow color. The Heisenberg exchange interaction parameters $J_{1}^{+}$, $J_{1}^{-}$, and $J_{2}$ are highlighted for selected atoms, with thin red arrows showing the atom pairs involved in these interactions. Rendered using VESTA [128].
Material Method $J_{1}^{+}$ $J_{1}^{-}$ $J_{2}$ $J_{3}^{+}$ $J_{3}^{-}$ $J_{4}$ $J_{5}$
NiO MFT -0.12 -0.16 13.88 0.01 0.00 0.09 0.03
$\Delta E$ -1.21 -0.95 12.57 -0.04 -0.04 0.49 0.01
MnO MFT 1.23 0.83 1.03 0.00 0.01 0.03 0.02
$\Delta E$ 0.47 0.36 0.59 0.00 0.00 0.05 0.01
@@
Due to the small rhombohedral distortion of the lattice at low temperatures, certain exchange interactions become inequivalent for parallel and antiparallel spin orientations. In particular, the nearest-neighbor exchange interaction $J_{1}$ splits into two distinct parameters, $J_{1}^{-}$ and $J_{1}^{+}$ – a notation introduced by Lines and Jones [129] – to distinguish between interactions involving parallel and antiparallel spin pairs, respectively. This splitting originates from magnetoelastic coupling induced by the distortion. In the absence of such structural distortions, these parameters are equal, i.e., $J_{1}^{-}=J_{1}^{+}$. In contrast, the next-nearest-neighbor exchange interaction $J_{2}$ occurs only between antiparallel spin pairs, and thus a single parameter is sufficient to describe it. These exchange interactions are illustrated in Fig. 1; interactions beyond the second-neighbor shell are not shown. A detailed analysis of the physical origin and behavior of exchange interactions in these systems can be found, for example, in Refs. 24, 25.
@@
III.2 First-principles calculations of exchange parameters using the direct MFT and $\Delta E$ methods
@@
As discussed, we first compute magnetic exchange parameters using two direct first-principles approaches: MFT and $\Delta E$. Both methods are applied consistently on top of the same DFT+$U$ ground state, using identical Hubbard $U$ values obtained from DFPT. The inclusion of $U$ is essential for NiO and MnO, as it corrects self-interaction errors of (semi-)local functionals and thus properly localizes the $d$ electrons. This consistent setup contrasts with previous studies [25, 38, 35, 105], where a single method was used to compute exchange parameters but with varying ground-state descriptions. By fixing the ground state using DFT+$U$, we investigate the accuracy of each method (MFT and $\Delta E$) in predicting the MIPs, similar to the strategy in Ref. 24 (which employed the SIC method for the ground-state description). Moreover, using first-principles $U$ values removes ambiguity in the choice of $U$, avoiding additional spread in the computed MIPs.
@@
Each of the methods considered yields individual exchange parameters $J$, which must be computed for as many neighbor shells as needed to ensure convergence of the magnon dispersions. For both NiO and MnO, we find that including exchange interactions up to the fifth nearest neighbor shell is necessary to achieve stable (i.e., positive-definite) magnon dispersions with convergence within a few meV. The corresponding values are listed in Table 1. While most previous studies [104, 24, 25, 38, 35, 105, 106, 107] limit the analysis to only $J_{1}$ and $J_{2}$ (often neglecting the splitting of $J_{1}$), our results highlight the importance of computing further-neighbor couplings to obtain stable and quantitatively accurate magnon dispersions. We also find that the splitting between parallel and antiparallel spin interactions occurs not only for the first nearest-neighbor shell ($J_{1}$) but also for the third nearest-neighbor shell ($J_{3}$) when the MFT method is used. All reported $J$ values are expressed using spin quantum numbers $S=1$ for Ni2+ and $S=5/2$ for Mn2+ (see Eq. (1)). Alternative definitions, such as setting $S$ equal to half the calculated magnetic moment, exist in the literature, but suffer from ambiguity in the definition of the magnetic moment. Therefore, the choice of $S$ must be taken into account when comparing the exchange parameters across different computational studies or with experimentally fitted values, as will be discussed below.
@@
Figure 2: Magnon dispersions for (a) NiO and (b) MnO computed using LSWT with Heisenberg exchange parameters from MFT and $\Delta E$ are shown as blue and red solid lines, respectively (see Table 1 and Eq. (1)). The TDDFPT+$U$ magnons dispersions are computed directly from the dynamical spin susceptibility and are shown as green dots. Experimental magnon dispersions, shown as black dots, are taken from inelastic neutron scattering measurements reported in Refs. 130, 131. The green and black dashed lines represent LSWT fits to the TDDFPT+$U$ and experimental data, respectively, based on the effective Heisenberg Hamiltonian of Eq. (2). All theoretical results are based on the same DFT+$U$ ground state.
@@
For NiO, both MFT and $\Delta E$ yield similar values for the dominant antiferromagnetic exchange interaction $J_{2}$, with MFT predicting a value approximately 10% higher than that from $\Delta E$. Consistently with previous computational studies, $J_{2}$ is the largest exchange parameter. In contrast, $J_{1}^{+}$ and $J_{1}^{-}$ are 1–2 orders of magnitude smaller and have opposite sign to $J_{2}$. However, their values differ significantly between methods: MFT predicts $|J_{1}^{+}|<|J_{1}^{-}|$, which favors the stabilization of the observed antiferromagnetic spin structure. In contrast, $\Delta E$ gives $|J_{1}^{+}|>|J_{1}^{-}|$, which destabilizes the magnon spectrum unless further-neighbor interactions are included. For the third to fifth neighbor interactions, both methods predict very small values. MFT yields negligible values, with only $J_{4}$ being about 1.5–2 times smaller than $J_{1}^{+}$ and $J_{1}^{-}$. A similar trend is observed in $\Delta E$, where only $J_{4}$ is sizable. Although these couplings are minor compared to $J_{2}$, they are essential in the $\Delta E$ approach to stabilize magnons, while their impact in MFT is minimal (see Fig. S1(a) in the SM [115]).
@@
For MnO, the dominant exchange interactions are $J_{1}^{+}$, $J_{1}^{-}$, and $J_{2}$, all having similar magnitudes and the same (positive) sign within each method, unlike in NiO. However, MFT and $\Delta E$ predict different trends: MFT gives $|J_{1}^{+}|>|J_{2}|>|J_{1}^{-}|$, while $\Delta E$ yields $|J_{2}|>|J_{1}^{+}|>|J_{1}^{-}|$. The inclusion of these three parameters alone is sufficient to stabilize magnons in both approaches. However, their absolute values are consistently larger in MFT than in $\Delta E$ by roughly a factor of two, leading to similarly scaled differences in magnon energies [see Fig. S1(b) in the SM [115]]. Third to fifth nearest-neighbor interactions are 1–2 orders of magnitude smaller than $J_{1}^{+}$, $J_{1}^{-}$, and $J_{2}$. Despite their small magnitude, including these further-neighbor interactions is necessary to achieve convergence of the magnon dispersion. Omitting them leads to energy deviations of up to 4 meV for certain magnon momenta, corresponding to errors of about 15%. Thus, while these interactions are not essential for magnon stability (as they are in NiO) they are important for the quantitative comparison with the experimental data.
III.3 Magnon dispersions
@@
We now compare the three computational approaches considered, to predict magnon dispersions in NiO and MnO. The first two are based on LSWT using Heisenberg exchange parameters $J$ obtained from MFT and $\Delta E$, as discussed in the previous section. The third approach computes magnon dispersions directly via TDDFPT+$U$ by evaluating the dynamical spin susceptibility [39]. A brief overview of each method is provided in Sec. S3 of the SM [115]. All methods are applied consistently on top of the same DFT+$U$ ground state, enabling a direct comparison focused on the physical approximations inherent to each method in describing magnetic excitations. The computed magnon spectra are benchmarked against inelastic neutron scattering data from Refs. 130, 131.
@@
The primitive magnetic unit cell of NiO and MnO contains two antiferromagnetically aligned magnetic ions due to the AFII spin configuration. In the absence of magnetic anisotropy, this leads to two degenerate magnon modes. However, weak magnetocrystalline anisotropies (arising from spin-orbit coupling) lift this degeneracy near the $\Gamma$ point, resulting in a small magnon gap at $\Gamma$ of less than 5 meV in both materials [129, 130, 131]. Although this gap is small in NiO and more pronounced in MnO, relative to the maximum magnon energies in each material, we follow common practice in the literature and neglect magnetocrystalline anisotropy, treating the magnon gap at $\Gamma$ as zero. This simplification is justified because our primary aim is to compare the accuracy of different computational methods in predicting magnon dispersions across the full BZ, rather than focusing on fine features near the $\Gamma$ point.
@@
Figure 2 shows the magnon dispersions computed using the three methods along the $\Gamma$–M and $\Gamma$–X directions in the BZ for rhombohedrally distorted NiO and MnO. TDDFPT+$U$ and LSWT based on $\Delta E$-derived exchange parameters both reproduce the experimental data well, capturing the overall trends and curvatures of the magnon bands. The maximum magnon energies are overestimated on average by about 15 meV (14%) for NiO and 3 meV (17%) for MnO. Between the two methods, LSWT with $\Delta E$-derived parameters shows better agreement with experiment than TDDFPT+$U$. This good agreement of both methods with experiments is remarkable given that both methods are fully first-principles and rely exclusively on ab initio Hubbard $U$ values, without any empirical fitting. In contrast, LSWT using MFT-derived $J$ values, while qualitatively reproducing the general features of the dispersion, significantly overestimates the magnon energies. The deviations reach 46 meV (40%) for NiO and 25 meV (139%) for MnO. Such a large deviation of magnon energies from MFT, especially for MnO, is quite striking when compared to the other two methods.
@@
A key feature of the magnon dispersion in NiO and MnO is the finite magnon energy at the M point in the BZ. This feature arises only when the rhombohedral lattice distortion is taken into account [35, 39], which splits the nearest-neighbor (and some further-neighbor) exchange interaction into two distinct parameters, $J_{1}^{+}$ and $J_{1}^{-}$. If the distortion is neglected, then $J_{1}^{+}=J_{1}^{-}$, and the magnon energy at M becomes zero. In our calculations, which include the rhombohedral distortion, all three methods correctly predict a finite magnon energy at M. However, the values obtained from TDDFPT+$U$ and LSWT based on $\Delta E$-derived exchange parameters agree much better with experiment than those from LSWT based on MFT-derived parameters. This magnon energy at M is relatively small in NiO compared to its maximum magnon energy, but significantly larger in MnO. This difference reflects the relative magnitudes of the exchange interactions: in NiO, $J_{2}$ dominates and exceeds $J_{1}^{+}$ and $J_{1}^{-}$ by 1–2 orders of magnitude, whereas in MnO, all three parameters are of similar magnitude (see Table 1).
@@
Figure 3: Effective exchange parameters for NiO within the minimal Heisenberg model [Eq. (2)]: (a) $\mathcal{J}_{1}^{+}$, (b) $\mathcal{J}_{1}^{-}$, and (c) $\mathcal{J}_{2}$, all in meV, extracted by fitting magnon dispersions using the analytical LSWT expression (see Sec. S3 in the SM [115]) for four data sets: $(i)$ LSWT dispersions with MFT-derived $J_{1}$$J_{5}$ values, $(ii)$ LSWT dispersions with $\Delta E$-derived $J_{1}$$J_{5}$ values, $(iii)$ TDDFPT+$U$ magnon dispersions from the dynamical spin susceptibility, $(iv)$ Experimental data from Ref. 130. Theoretical results are shown as histograms, while experimental values are indicated by horizontal dashed lines. All theoretical results are based on the same DFT+$U$ ground state.
@@
Figure 4: Effective exchange parameters for MnO within the minimal Heisenberg model [Eq. (2)]: (a) $\mathcal{J}_{1}^{+}$, (b) $\mathcal{J}_{1}^{-}$, and (c) $\mathcal{J}_{2}$, all in meV, extracted by fitting magnon dispersions using the analytical LSWT expression (see Sec. S3 in the SM [115]) for four data sets: $(i)$ LSWT dispersions with MFT-derived $J_{1}$$J_{5}$ values, $(ii)$ LSWT dispersions with $\Delta E$-derived $J_{1}$$J_{5}$ values, $(iii)$ TDDFPT+$U$ magnon dispersions from the dynamical spin susceptibility, $(iv)$ Experimental data from Ref. 131. Theoretical results are shown as histograms, while experimental values are indicated by horizontal dashed lines. All theoretical results are based on the same DFT+$U$ ground state.
III.4 Fitting effective exchange parameters of the minimal Heisenberg model